OpenStudy (anonymous):

Find if a = 34, b = 28 and c = 15. Law of cosines

OpenStudy (campbell_st):

so which angle to you need to find...?

OpenStudy (anonymous):

Oh sorry, measure of angle A

OpenStudy (campbell_st):

so the law of cosines for angle A is \[\cos(A) = \frac{b^2 + c^2 - a^2}{2bc}\] that will give a decimal value less than 1 then to find A its \[A = \cos^{-1}(Answer)\] hope it helps

OpenStudy (anonymous):

It did, thanks!